OpenStudy (anonymous):

x-2y >-8 __ 7x+y< -6

OpenStudy (anonymous):

the - goes under the symbol under the first greater than sign

OpenStudy (anonymous):

Hmm... I've actually forgotten how to solve this... I know what the answer is, but let's see if I can actually solve it.

OpenStudy (anonymous):

x-2y >= -8 -x+2y <= 8 2y <= x + 8 y <= 1/2 x + 4 7x+y< -6 -7x - y > 6 -7x - 6 > y I'm not sure what to do from here...

OpenStudy (anonymous):

If you graph the two, you'll see what the answers are. (They're x >= -4/3 and x< -4/3) SOmeone else please explain this.